1/2(6a+4)+1/2(6a-4)
1
Expert's answer
2012-03-02T07:01:32-0500
1/{2(6a+4)}+1/{2(6a-4)}=1/2{(6a-4+6a+4)/((6a+4)(6a-4))}=1/2*12a/(36a2-16)=6a/4/(9a2-4)=3a/{2(9a2-4)}